In a music poll of 25 students. 15 liked rock, 9 liked jazz, and 12 liked rap. 4 students liked all 3 types of music. 4 liked only rock and jazz, and 5 liked only rock and rap. If no students liked only jazz and rap, how many students did not like any of these types of music?
Answer by stanbon(75887)

In a music poll of 25 students. 15 liked rock, 9 liked jazz, and 12 liked rap. 4 students liked all 3 types of music. 4 liked only rock and jazz, and 5 liked only rock and rap. If no students liked only jazz and rap, how many students did not like any of these types of music?
---------------
Sketch a rectangle with 3 intersecting circles labeled rock, jazz, rap
----
Put 4 at the intersection of all three circles
Put 4 at the intersection of rock and jazz
Put 5 at the intersection of rap and rock
Put 0 at the intersection of jazz and rap
----------------
Total in the circles:: 4 + 4 + 5 = 13
Ans: # none of the circles = 25-13 = 12
-------------------
Cheers,
Stan H.
---------------